Problem 13-4 The current aggregate demand requirements for a firm are shown below for the next six months MonthMay June July Aug Sept Oct Demand 260 240 240 240 270 290 The firm always plans to meet all demand. The firm currently has 260 workers capable of producing 260 units in a month (1 unit/worker). The workforce can be increased (at a cost of $650 (at a cost of $1,300 per worker). Inventory holding cost is $150 per unit per month. The firm currently has 40 units of inventory on hand, and it would like to have 40 units available at the e production cost is $3,700 per unit. Assume hiring and layoff/firing, if necessary, occur at the beginning of the month a. What should the aggregate plan be if the inventory holding cost is to be minimized? (Leave no cells blank be certain to enter wherever required.) Regular Production Ending Invento Number of Workers Month May June July August September October Total Demand Hire Fire 260 240 240 240 270 290 1,540 260 240 240 240 270 290 1,540
Answer #1

Since the firm wants to meet demand, so hiring and firing will happen every month according to demand.

The costs would be Production cost, Hiring cost, Layoff cost and Inventory cost

The problem in excel looks like as below

Aggregate 3 Hiring cost 650pe per worker per worker per unit per m per unit 4 Layoff cost Inventory Cost 150 6 Production Cost 3700 8 Month Beg Inv elar Pr End Inv Number of woker: Hire Production Cost Inventory cost Hiring cost Layoff Cost Total Cost 10 May 11 June 12 July 13 August 14 September 15 October 40 E10 -E11 -E12 240 -E13 -E14 290 D10* $B$6 =MAX(F11-F12,0) =012*SBS6 -MAX(F13-F14,0) D14 SB$6 260 260 240 812+D12-C121 240 240 240 240 =MAX(F12-F11,0) -E12 $B$5 813+D13-C13 240 240 270 B14+D14-c14 270 -MAX(F14-F13,0) -E14 SB$5 -MAX(F15-F14,0) -MAX(F14-F15,0) D15*SB$6 -SUM M10:M16) 17 Total 18Thus the output is as below